OpenStudy (anonymous):

Help! solve this equation for x (2/(x+5)+(1/(x-5)=(16/(x^2-25)

OpenStudy (anonymous):

\[\frac{2}{x+5}+\frac{1}{x-5}=\frac{16}{x^2-25}\] \[\frac{2(x-5)+1(x+5)}{x^2-25}=\frac{16}{x^2-25}\] \[2x-10+x+5=\frac{16}{\cancel{(x^2-25)}}*\cancel{(x^2-25)}\] \[3x-5=16\] \[3x=16+5\] \[x=\frac{21}{3}=7\]